OpenStudy (anonymous):

Find the derivative of the trigonometric function. g(t)=quadroot(t)+6csct

OpenStudy (anonymous):

\[g(x)=\sqrt[4]{t}-6csct\]

OpenStudy (anonymous):

t^1/4-6csct

OpenStudy (anonymous):

1/4t^-3/4 - 6csc(t)cot(t)

OpenStudy (anonymous):

1/4t^3/4 - 6 csc(t)cot(t) ????

OpenStudy (anonymous):

\[1/4t ^{2/3} - 6\csc(t)\cot(t)\]

OpenStudy (anonymous):

oops that supposed to be 3/4 >.<

OpenStudy (anonymous):

\[1/4t ^{3/4} - 6\csc(t)\cot(t)\]

OpenStudy (anonymous):

thanks a lot

OpenStudy (mimi_x3):

wait..

OpenStudy (mimi_x3):

\[\huge \frac{1}{4} t^{\frac{-3}{4}} \]

OpenStudy (anonymous):

i think when i drop the t to the bottom i lose the - sign

OpenStudy (anonymous):

on the exponent

OpenStudy (anonymous):

1/(4t^3/4) - 6 csc(x)cot(x)

OpenStudy (anonymous):

maybe thats a better way to write it

OpenStudy (mimi_x3):

\[\huge t^{\frac{1}{4}} =>\frac{1}{4} t^{\frac{1}{4}-1} =>\frac{1}{4} t^{-\frac{3}{4}} \]

OpenStudy (mimi_x3):

Yeah, that is a better way of writing it, it looks nicer.

OpenStudy (anonymous):

thanks a lot mimi

OpenStudy (anonymous):

you are a life saver :)

OpenStudy (mimi_x3):

yw (: